If a dice Is thrown 2 times then what is the probability of not getting 5 everytime

Grade:10

1 Answers

Arun
6 years ago
Dear Rajat
 
Thus we want that in first throw, 5 should not come and also in second throw it should not come
 
Hence probability = 5/6 *5/6 = 25/36
 
 
Regards
Arun (askIITians forum expert)